Exercise 8-1 Schedule of Expected Cash Collections [LO8-2]

Silver Company makes a product that is very popular as a Mother’s Day gift. Thus, peak sales occur in May of each year, as shown in the company’s sales budget for the second quarter given below:

April May June Total
Budgeted sales (all on account) $410,000 $610,000 $210,000 $1,230,000

From past experience, the company has learned that 25% of a month’s sales are collected in the month of sale, another 60% are collected in the month following sale, and the remaining 15% are collected in the second month following sale. Bad debts are negligible and can be ignored. February sales totaled $340,000, and March sales totaled $370,000.

Required:

1. Prepare a schedule of expected cash collections from sales, by month and in total, for the second quarter.

2. What is the accounts receivable balance on June 30th?
Answer #1

Requirement 1

Schedule of Expected Cash Collections
April May June Total
February Sales $ 51,000 $ 51,000
March Sales $ 222,000 $ 55,500 $ 277,500
April Sales $ 102,500 $ 246,000 $ 61,500 $ 410,000
May Sales $ 152,500 $ 366,000 $ 518,500
June Sales $ 52,500 $ 52,500
Total Cash Collection $ 375,500 $ 454,000 $ 480,000 $ 1,309,500

Requirement 2

Silver company
Accounts receivables at june 30
May sales $ 91,500
June sales $ 157,500
Total Accounts receivables at june 30 $ 249,000

Working

Schedule of Expected Cash Collections
April May June
February Sales =340000*15%
March Sales =370000*60% =370000*15%
April Sales =410000*25% =410000*60% =410000*15%
May Sales =610000*25% =610000*60%
June Sales =210000*25%

.

Accounts Receivables Balance
Balance of may sales (15% of sale) $ 91,500
Balance of June Sale (75% of sale) $ 157,500
Total Receivables at June 30 $ 249,000